# SSRLBot: Designing and Developing an LLM-based Agent using Socially Shared Regulated Learning

### v1

@@ Abstract
Large language model (LLM)-based agents are increasingly used to support human experts by streamlining complex tasks and offering actionable insights. However, their application in multi-professional decision-making, particularly in teamwork contexts, remains underexplored. This design-based study addresses that gap by developing LLM functions to enhance collaboration, grounded in the Socially Shared Regulation of Learning (SSRL) framework and applied to medical diagnostic teamwork. SSRL emphasizes metacognitive, cognitive, motivational, and emotional processes in shared learning, focusing on how teams manage these processes to improve decision-making.
This paper introduces SSRLBot, a prototype chatbot designed to help team members reflect on both their diagnostic performance and key SSRL skills. Its core functions include summarizing dialogues, analyzing SSRL behaviors, evaluating diagnostic outcomes, annotating SSRL markers in conversation, assessing their impact on performance, and identifying interpersonal regulatory dynamics. We compare SSRLBot’s capabilities with those of Gemini-1.5, GPT-3.5, and Deepseek-R1 in a case study. SSRLBot demonstrates stronger alignment with SSRL theory, offering detailed evaluations that link behaviors to regulatory dimensions and suggesting improvements for collaboration. By integrating SSRL theory with LLM capabilities, SSRLBot contributes a novel tool for enhancing team-based decision-making and collaborative learning in high-stakes environments, such as medical education.
@@
Generative Artificial Intelligence Large Language Models Socially shared regulation Human-Centred Design Medical Diagnosis.
@@ 1 Introduction
Team actions—decisions and operations carried out collaboratively—are essential for achieving success. However, without the joint efforts of team members in adopting appropriate strategies, operations, and regulations, team action alone may not guarantee success [14]. This is especially true in high-stakes decision-making contexts, such as clinical reasoning, where a single decision can lead to significantly different outcomes [5]. The effectiveness of team decision-making is closely tied to Socially Shared Regulation of Learning (SSRL, [12]), a framework that emphasizes the collaborative regulation of learning processes as key to improving decision-making outcomes. In medical teamwork, team members’ interactions through conversation plays a pivotal role in their diagnostic outcome, as they could either direct to the correct or an opposite result [2]. Medical simulation has been pivotal for medical professionals to practice diagnostic skills, and it paves the way for understanding how teamwork works in this field [3, 13].
The purpose of this design-based study is to design and develop a large language model (LLM) based agent - SSRLBot, for teamwork evaluation using case conversations in a medical diagnostic task. Grounded from the theory of socially shared regulated learning (SSRL), SSRLBot enables team members to reflect on their diagnostic performance as a team and the key SSRL skills being used to foster team diagnostic success in medical simulation. It also generates instant reports for medical supervisors, providing insights into practitioners’ performance and collaboration quality, paving the way for customized scaffolding.
@@ 2.1 SSRL in Team Work
Socially shared regulation is essential in teamwork. It involves team members’ mutual efforts in regulating team performance through effective interactions. Socially shared regulated learning reflects learning strategies in educational contexts where two or more learners are involved. Studies in collaborative learning have found that groups that perform better show more SSRL strategies in teamwork [9]. However, the mechanism of SSRL is complex and hard to be captured, as dynamic interpersonal interactions occur where team members are influencing each other. SSRL guides interactions in teamwork from four dimensions, namely meta-cognitive, socio-cognitive, socio-motivational, and socio-emotional aspects. The meta-cognitive aspect of interactions (short in meta-cognitive interaction in the following content) refers to conversations about about the team’s cognitive progress [6, 11]. Cognitive interaction include conversations about the team’s actual cognitive activities, such as planning, performing, and reflecting about the task and team performance [11]. Meta-cognitive interaction to teamwork is like an umbrella to an outdoor walking with weather forecast one, it prevents rains when necessary and can be folded when the sun comes out. It functions as a signal for cognitive actions and monitor important cognitive interactions in teamwork. Socio-emotional interaction, or social emotional interaction is efforts team members made to maintain cohesive and respectful social interaction through team dynamics [8]. Healthy and positive socio-emotional interaction enables team members to engage more effectively in cognitive interaction and lead to better team outcome. Socio-motivational interaction is often discusses with socio-emotional interactions [10, 8], as it is pivotal to motivate team spirits and triggers team members to stay tuned and engaged in teamwork.
SSRL research often conducts discourse analysis in computer-supported collaborative learning (CSCL) contexts, as it allows for the identification of learners’ situational social, cognitive, emotional, and motivational states through transcriptions of their dialogues or written texts [7]. Integrating trace data from think-aloud protocols and CSCL logfiles represents a significant advancement in capturing learners’ thought processes and actions, providing a more objective perspective on SSRL [1]. Most studies analyze verbal data through human coding, which, despite expert evaluation, remains labor-intensive and prone to bias. Large language models (LLMs) learn word usage patterns and apply them to natural language processing tasks [18], offering a more efficient alternative that often matches or surpasses human performance in content understanding, idea generation, and decision-making.
@@ 2.2 LLM Agent in Education
LLM agents have recently debuted in educational contexts. For example, PBChat is designed to identify and provide solutions on students’ problem behaviors using teacher-parent concersations. LLMAgent-CK [20] is designed as a multi-agent framework to identify middle-school teachers’ mathematical content knowledge learning goals from their responses. Other LLM-based agents are designed as virtual tutor or assistants in diverse subjects (e.g., Ruffle & Riley: Biology [15]; Jill Watsons, course readings [17]). Furthermore, other LLM-based agents have been designed for advancing educational research (e.g., Vizchat for visualizing multimodal learning analytics [19]).
Despite previous applications, their potential for medical education remains unexplored. The SSRL theory-driven design could provide valuable insights into teamwork evaluation. Notably, SSRLBot is the first LLM-based agent designed to assess medical teams’ SSRL interactions and their impact on diagnostic performance.
@@ 3.2 Functional Realization
The instruction description consists of a preamble and function descriptions. The preamble’s setting helps SSRLBot clearly understand its role and continue playing this role in a scene for dialogue. We set the SSRLBot as a professional researcher with enough SSRL knowledge and health knowledge, be able to analyze and evaluate data. Moreover, we developed three core functions in SSRLBot: (1) evaluating each team member’s interpersonal influence on others, (2) annotating each member’s SSRL skills, and (3) providing suggestions for growth.
3.3 Tuning
@@ 4 Illustrative Case Study
4.1 Materials
This section presents a case study using SSRLBot to generate contextualized explanations within the SSRL framework. With ethics approval, we analyzed dialogues from a dyad of medical residents at a large North American university engaged in a team diagnostic task. Figure 2 illustrates functions of SSRLBot.
Figure 2: The Demonstration of SSRLBot Functions.
Specifically, by analyzing the given sample, SSRLBot demonstrates its three core functions. First, it identifies each team member’s interpersonal SSRL influence by linking dialogue evidence and summarizing their roles, as shown in Function 1 on the left of the figure. Second, it locates each conversation line and visualizes a comparison of SSRL skills between the two speakers, as seen in the top right corner (Function 2). Third, SSRLBot evaluates the team conversation based on each member’s SSRL skills and suggests areas for improvement (Function 3). Together, these functions enable SSRLBot to provide actionable insights into team dynamics, fostering more effective collaborative learning and problem-solving.
4.2 Contextualized Evaluation
For contextualized evaluation, we compare SSRLBot with other LLMs, including Gemini-1.5, ChatGPT-3.5, and Deepseek-R1. The comparisons are guided by: 1. contextualized description (providing customized details for each human involved in the conversation) and 2. theoretical adherence (alignment with SSRL theory).
All selected LLMs demonstrated their ability to understand the conversation’s goal, identify key information, and generate a general report based on the given content. However, ChatGPT 3.5 relied less on contextualized descriptions for each participant, limiting its usefulness for individuals aiming to improve their teamwork skills. Deepseek-R1 provided a more thorough evaluation by blending positive and critical feedback on participants’ SSRL skills and their impact on diagnostic performance. However, its theoretical alignment could be improved by explicitly linking behaviors to SSRL dimensions and offering more personalized feedback. Gemini-1.5 provided contextualized descriptions; for example, it acknowledged humor’s role within the team and aligned with SSRL theory. However, its analysis lacked actionable feedback for human-centered reflection. In contrast, SSRLBot delivered a more detailed evaluation by categorizing SSRL skills within the metacognitive, cognitive, emotional, and motivational dimensions while highlighting specific conversational examples. Additionally, it provided a comprehensive analysis by offering customized recommendations for each participant, making it clearer where teamwork improvements were needed. By linking evidence from the conversation to each participant’s SSRL skills, SSRLBot offered a clearer blueprint for enhancing teamwork, effectively guiding users on which specific dimensions to adapt for improved team decision-making.
@@ 5 Conclusion and Future Work
In this paper, we introduced SSRLBot, a knowledge-based LLM agent for teamwork conversation evaluation, grounded from the socially shared regulated learning framework. Integrating the psychological concepts of meta-cognitive, socio-cognitive, socio-emotional, and socio-motivational interactions, SSRLBot aims to facilitate customized feedback for communicators through evaluating their conversations in teamwork contexts. This theory-driven knowledge empowering design enhances LLM functions that support teamwork enhancement through customized feedback for individuals and groups. Our case study showcases SSRLBot’s outstanding capabilities: it accesses to every loop of conversation for a comprehensive summary, differentiates each team members’ SSRL capacity and their interpersonal influence on each other, visualizes SSRL components with contextually relevant explanations, and offers detailed insights for future growth for teamwork performance. These functionalities can be beneficial for team leaders, educational stakeholders, and team members who want to excel in teamwork tasks through effective communications or interactions.
Future research can enhance SSRLBot by expanding its application beyond dyadic medical diagnosis to diverse contexts, enabling a more comprehensive evaluation of its theory-driven functions. Additionally, while researchers assessed its contextual functions, incorporating participant feedback would provide a more robust measure of its reliability and informativeness.

### v2

# SSRLBot: Designing and Developing a Large Language Model-based Agent using Socially Shared Regulated Learning

@@ Abstract
Large language model (LLM)–based agents have emerged as pivotal tools in assisting human experts across various fields by transforming complex tasks into more efficient workflows and providing actionable stakeholder insights. Despite their potential, the application of LLM-based agents for medical education remains underexplored. The study aims to assist in evaluating the students’ process and outcomes on medical case diagnosis and discussion while incorporating the theoretical framework of Socially Shared Regulation of Learning (SSRL) to assess student performance. SSRL emphasizes metacognitive, cognitive, motivational, and emotional interactions, highlighting the collaborative management of learning processes to improve decision-making outcomes. Grounded in SSRL theory, this tool paper introduces SSRLBot, an LLM-based agent designed to enable team members to reflect on their diagnostic performance and the key SSRL skills that foster team success. SSRLBot’s core functions include summarizing dialogue content, analyzing participants’ SSRL skills, and evaluating students’ diagnostic results. Meanwhile, we evaluated SSRLBot through diagnostic conversation data collected from six groups (12 participants, 1926 conversational turns). Results showed that SSRLBot can deliver detailed, theory-aligned evaluations, link specific behaviors to SSRL dimensions, and offer actionable recommendations for improving teamwork. The findings address a critical gap in medical education, advancing the application of LLM agents to enhance team-based decision-making and collaboration in high-stakes environments.
@@
Generative Artificial Intelligence Large Language Models Socially Shared Regulation Medical Education Conversational Agents.
@@ 1 Introduction
Team actions—decisions and operations carried out collaboratively—are essential for achieving success. However, team action alone may not guarantee success without the joint efforts of team members in adopting appropriate strategies, operations, and regulations  [15]. This is especially true in high-stakes decision-making contexts, such as clinical reasoning, where a single decision can lead to significantly different outcomes  [5]. The effectiveness of team decision-making is closely tied to the Socially Shared Regulation of Learning (SSRL,  [13]). This framework emphasizes the collaborative regulation of learning processes as key to improving decision-making outcomes. In medical teamwork, team members’ interactions through conversation play a pivotal role in their diagnostic outcome, as they could either direct to the correct or an opposite result  [2]. Medical simulation has been pivotal for medical professionals to practice diagnostic skills and paves the way for understanding how teamwork works in this field  [3, 14].
In this study, we designed and developed a large language model (LLM)-based agent, SSRLBot. SSRLBot can evaluate students’ performance in medical case diagnostics and conversations, judge the diagnosis results, and identify students’ SSRL skills in their conversations. Grounded on the theory of SSRL, SSRLBot enables students to reflect on their diagnostic performance as a team and the key SSRL skills used to foster team diagnostic success in medical simulation. Additionally, we used six groups’ diagnostic conversation data, including 12 participants and 1926 conversational turns, to evaluate this LLM-based agent.
@@ 2.1 SSRL in Team Work
Socially shared regulation is essential in teamwork. It involves team members’ mutual efforts in regulating team performance through effective interactions. Socially shared regulated learning reflects learning strategies in educational contexts where two or more learners are involved. Studies in collaborative learning have found that groups that perform better show more SSRL strategies in teamwork  [10]. However, the mechanism of SSRL is complex and hard to be captured, as dynamic interpersonal interactions occur where team members are influencing each other. SSRL guides interactions in teamwork from four dimensions, namely meta-cognitive, socio-cognitive, socio-motivational, and socio-emotional aspects. The meta-cognitive aspect of interactions (short in meta-cognitive interaction in the following content) refers to conversations about the team’s cognitive progress  [6, 12]. Cognitive interaction includes conversations about the team’s actual cognitive activities, such as planning, performing, and reflecting on the task and team performance  [12]. Meta-cognitive interaction to teamwork is like an umbrella to an outdoor walking with a weather forecast; it prevents rains when necessary and can be folded when the sun comes out. Socio-emotional interaction, or social-emotional interaction, is team members’ efforts to maintain cohesive and respectful social interaction through team dynamics  [8]. Healthy and positive socio-emotional interaction enables team members to engage more effectively in cognitive interaction and leads to better team outcomes. Socio-motivational interaction is often discussed with socio-emotional interactions  [11, 8], as it is pivotal to motivating team spirits and triggers team members to stay tuned and engaged in teamwork.
SSRL research often conducts discourse analysis in computer-supported collaborative learning (CSCL) contexts, as it allows for the identification of learners’ situational social, cognitive, emotional, and motivational states through transcriptions of their dialogues or written texts [7]. Integrating trace data from think-aloud protocols and CSCL logfiles represents a significant advancement in capturing learners’ thought processes and actions, providing a more objective perspective on SSRL [1]. Most studies analyze verbal data through human coding, which, despite expert evaluation, remains labor-intensive and prone to bias. LLMs learn word usage patterns and apply them to natural language processing tasks [19], offering a more efficient alternative that often matches or surpasses human performance in content understanding, idea generation, and decision-making.
@@ 2.2 LLM Agent in Education
LLM agents have recently debuted in educational contexts. For example, PBChat is designed to identify and provide solutions to students’ problem behaviors using teacher-parent conversations. LLMAgent-CK  [21] is designed as a multi-agent framework to identify middle-school teachers’ mathematical content knowledge learning goals from their responses. Other LLM-based agents are designed as virtual tutors or assistants in diverse subjects (e.g., Ruffle & Riley: Biology  [16]; Jill Watsons, course readings  [18]). Furthermore, other LLM-based agents have been designed for advancing educational research (e.g., Vizchat for visualizing multimodal learning analytics  [20]). Despite previous applications, their potential for medical education remains unexplored. The SSRL theory-driven design could provide valuable insights into teamwork evaluation. Notably, SSRLBot is the first LLM-based agent designed to assess medical teams’ SSRL interactions and their impact on diagnostic performance.
@@ 3.2 Functional Realization
The instruction description consists of a preamble and function descriptions. The preamble’s setting helps SSRLBot clearly understand its role and continue playing this role in a scene for dialogue. We set the SSRLBot as a professional researcher with enough SSRL knowledge and health knowledge, be able to analyze and evaluate data. Moreover, we developed three core functions in SSRLBot: (1) identify the participant’s interpersonal influence on each, (2) compare the SSRL skills, and (3) summarize and provide suggestions on participants. Figure 2 illustrates functions of SSRLBot. To enable SSRLBot to identify corresponding SSRL skills effectively and accurately, we developed an SSRL framework and integrated it into the system’s preamble. This framework consists of three SSRL skill layers. The first layer contains five core SSRL skills: meta-cognitive interaction, socio-cognitive interaction, socio-motivational interaction, socio-emotional interaction, and task execution. Each core SSRL skill also contains the second and third SSRL skill layers. Definitions and examples were provided for each skill to help the system make more accurate judgments.
Figure 2: The Demonstration of SSRLBot Functions.
@@ 4 Illustrative Case Study
In this section, we evaluated SSRLBot using six groups of authentic conversational data, which were collected from 12 medical residents recruited from a university in Montreal, Canada.
Data collection. Participants were paired up and required to discuss a medical case through the BioWorld system. BioWorld is a computer learning system that can provide a hospital simulation for students to discuss diseases using related knowledge  [9]. The duration of each conversation was around 30 minutes. The number of conversational turn in each group was between 123 and 600. We conducted basic data cleaning on the collected data. For example, we added a speaker (e.g., P1 for participant 1, P2 for participant 2, R for researcher) and the corresponding SSRL skill on each conversational turn. In addition, we evaluated the diagnostic results of each group with the standard diagnostic answer to determine whether each group’s results were correct.
Evaluation. Each group’s conversational data was saved as a case in a single file. In this study, we mainly evaluated the diagnostic performance of SSRLBot. We requested SSRLBot summarize the conversation content for each case and evaluate the diagnostic results. The responses provided by SSRLBot primarily include three sections: a summary of the conversation, an evaluation of diagnostic accuracy, and a conclusion. This agent enhanced the precision of its outcomes by extracting the keywords and critical conversational content as evidence. Researchers compared the SSRLBot’s results and the standard results and found that SSRLBot demonstrated 100% diagnostic accuracy.
Comparison. We mainly compare the performance of SSRLBot with other LLMs, including ChatGPT-3.5, Deepseek-R1, and Gemini-1.5. All selected LLMs demonstrated their ability to understand the conversation’s goal, identify key information, and generate a general report based on the given content. However, ChatGPT 3.5 relied less on contextualized descriptions for each participant, limiting its usefulness for individuals aiming to improve their teamwork skills. DeepSeek-R1 provided a more thorough evaluation by blending positive and critical feedback on participants’ SSRL skills and their impact on diagnostic performance. However, its theoretical alignment could be improved by explicitly linking behaviors to SSRL dimensions and offering more personalized feedback. Gemini-1.5 provided contextualized descriptions; for example, it acknowledged humor’s role within the team and aligned with SSRL theory. However, its analysis lacked actionable feedback for human-centered reflection. In contrast, SSRLBot can follow the SSRL framework to better identify the specific SSRL skills from the conversation while highlighting specific examples. Additionally, it provided a comprehensive analysis by linking evidence from the conversation to each participant’s SSRL skills, guiding researchers on which specific SSRL skill layers to adapt for improved team decision-making.
@@ 5 Conclusion and Future Work
This paper introduced SSRLBot, an LLM-based agent for medical education practice grounded in the SSRL framework. SSRLBot aims to facilitate customized feedback for learners through evaluating their conversations in the collaborative learning environment. This theory-driven knowledge-empowering design enhances LLM functions that support teamwork enhancement through customized feedback for individuals and groups. Our case study demonstrates that SSRLBot’s responses are highly accurate. The limitation of this study is the limited number of cases for evaluation. Although each case included extensive conversational data, it is important to conduct more cases to further evaluate the accuracy and effectiveness of this agent. Future research should expand the sample size and evaluate other functions of SSRLBot. In addition, future research can enhance SSRLBot by expanding its application beyond dyadic medical diagnosis to diverse contexts, enabling a more comprehensive evaluation of its theory-driven functions. Additionally, while researchers assessed its contex- tual functions, incorporating participant feedback would provide a more robust measure of its reliability and informativeness. Overall, the functionalities of this agent can be beneficial for medical education by helping assess learners’ performance and SSRL skills and providing effective summarization and evaluation evidence for researchers.